A will provides that Janice and Wanda should divide an estate in the ratio of 11 to 2. If Janice's share of estate is $56500, how much was the estate worth altogether? Round your answer to two decimal places.
Select the correct answer.
a. $66669.61
b. $66772.73
c. $66796.97
d. $66731.60

Answer by TimothyLamb(4379)

j/w = 11/2
j = 56500
---
j/w = 11/2
56500/w = 11/2
(2/11)56500 = w
w = 10272.73
---
j + w = 56500.00 + 10272.73 = $66772.73
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php